# Annabelle Michael Carrell    Albert Gong    Abhishek Shetty    Raaz Dwivedi    Lester Mackey

### v5

@@ C.2 Proof of kernel max seminorm bound 13
The triangle inequality and the covering property 136 together imply that, with probability $1$,
@@
for each $f\in\mathcal{F}$. Since $s\mapsto e^{ts}$ is increasing, the bound 141, the assumed sub-Gaussianity (Def. A.2), and the fact that $\mathcal{C}_{\epsilon,\mathcal{F}}$ belongs to $\mathbb{B}_{\mathcal{F}}$ imply that
@@ C.3 Proof of Lipschitz kernel max seminorm bound 15
Since $\mathbb{P}(\mathcal{E}^{c})\leq\delta/2$, it only remains to upper bound $\mathopen{}\mathclose{{}\left\|{(\mathbb{P}_{\textup{in}}-\mathbb{P}_{\textup{out}})\mathbf{k}}}\right\|_{\mathcal{Z},\mathcal{Z}}$ on $\mathcal{E}$ with probability at least $21-\delta^{\prime}/$ 2.
@@
Suppose $(X_{\theta})_{\theta\in\Theta}$ is a sub-Gaussian process with respect to $\rho$ on an event $\mathcal{E}$, and define the diameter $\mathop{\mathrm{missing}}{diam}(\Theta,\rho)\triangleq\sup_{\theta,\theta^{\prime}\in\Theta}\rho(\theta,\theta^{\prime})$, the covering number
@@
$\mathbb{P}_{\mathcal{E}}(\sup_{\theta,\theta^{\prime}\in\Theta}|X_{\theta}-X_{\theta^{\prime}}|\geq 8(\mathcal{J}(\Theta,\rho)+t))\leq 2\exp(-t^{2}/\mathop{\mathrm{missing}}{diam}(\Theta,\rho)^{2})\quad\text{for all}\quad t>0.$ (154)
@@
Consider the bounded-Hölder metric $\rho$ 152 for a kernel $\mathbf{k}$ on $\mathcal{X}=\mathbb{R}^{d}$ and a finite set $\mathcal{Z}\subset\mathcal{X}$. If $\mathbf{Z}$ is a matrix with one row corresponding to each element of $\mathcal{Z}$, $r=\mathrm{rank}({\mathbf{Z}})$, and $R=\max_{\boldsymbol{z}\in\mathcal{Z}}\|{\boldsymbol{z}}\|_{2}$, then, in the notation of Lem. C.5,
@@
$\mathop{\mathrm{missing}}{diam}(\mathcal{Z},\rho)$ $\leq D\triangleq\min(c,\nu\sqrt{\frac{32}{3}}\max_{\boldsymbol{z}\in\mathcal{Z}}\sqrt{\mathbf{k}(\boldsymbol{z},\boldsymbol{z})}),\quad\text{and}\quad$ (156)
@@
The diameter bound 156 follows directly from the definition of $\rho$ 152 and the fact $\max_{\boldsymbol{z},\boldsymbol{z}^{\prime}\in\mathcal{Z}}\|{\boldsymbol{z}-\boldsymbol{z}^{\prime}}\|_{2}\leq 2R$.
To establish the covering number bound 155, we let $\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$ be a compact singular value decomposition of $\mathbf{Z}$ so that
@@
so that $\mathbf{V}\mathcal{C}$ satisfies the criteria of 160. Since $|\mathbf{V}\mathcal{C}|\leq|\mathcal{C}|\leq(1+4R/\epsilon^{2})^{r}$ by Wainwright (2019, Lem. 5.2), we must also have $|\mathcal{C}_{\textup{ext}}|\leq(1+4R/\epsilon^{2})^{r}$.
Now, since $\mathcal{C}_{\textup{ext}}$ has minimum cardinality amongst sets satisfying 160, for each $\boldsymbol{z}^{\prime}\in\mathcal{C}_{\textup{ext}}$, there is some $\boldsymbol{z}\in\mathcal{Z}$ satisfying $\|{\boldsymbol{z}^{\prime}-\boldsymbol{z}}\|_{2}\leq\epsilon^{2}/2$ (or else $\boldsymbol{z}^{\prime}$ would be superfluous). Hence, there exists a set $\mathcal{C}_{\textup{int}}\subseteq\mathcal{Z}$ satisfying
@@
Hence, for $u=\frac{c\epsilon}{2\sqrt{R}}$, $\mathcal{N}(u;\mathcal{Z},\rho)\leq|\mathcal{C}_{\textup{int}}|\leq(1+c^{2}/u^{2})^{r}$. Since $\epsilon>0$ was arbitrary, we have established 155.
@@
on $\mathcal{E}$ with probability at least $21-\delta^{\prime}/$ 2. Combining this bound with the inequality 150 yields the result.
@@
Instantiate the notation of Algs. 1 and 175 and define the query set
@@
Together, Lem. F.3, the KMS quality bound of Thm. 1, and the KH-Compress$(0.5)$ sub-Gaussian constant $\nu$ of Prop. B.5 imply that, with probability at least $\frac{1}{2}$,
@@
$=\|{(\frac{1}{n}\mathbf{\widehat{D}})^{-1}}\|_{\max}\frac{1}{n}\|{\mathbf{A}\mathbf{1}_{n}-\widehat{\mathbf{A}}\mathbf{1}_{n}}\|_{\infty}\|{\mathbf{V}}\|_{\max}.$ (190)
@@
$\max_{j\in[n]}\|{\sum_{i=1}^{j}\boldsymbol{x}_{\pi_{k+1}(\pi_{k}^{-1}(i))}^{k}}\|_{2}\leq\frac{1}{2}\max_{j\in[n]}\|{\sum_{i=1}^{j}\boldsymbol{x}_{i}^{k}}\|_{2}+\frac{1}{2}\max_{j\in[n]}\|{\sum_{i=1}^{j}\epsilon_{i}^{k}\boldsymbol{x}_{i}^{k}}\|_{2}+\|{\sum_{i=1}^{n}\boldsymbol{x}_{i}^{k}}\|_{2}$ (202)
where $\pi_{k}^{-1}$ is the inverse permutation of $\pi_{k}$ and $\epsilon_{i}^{k}\triangleq 2(\mathbf{1}\mathopen{}\mathclose{{}\left[{\boldsymbol{x}_{i}^{k}\in\mathcal{X}_{\textup{out}}^{k}}}\right]-1)$.
@@
Fix any $j^{\star}\in\mathop{\mathrm{missing}}{argmax}_{j\in[n]}\|{\sum_{i=1}^{j}\boldsymbol{x}_{\pi_{k+1}(\pi_{k}^{-1}(i))}^{k}}\|_{2}$. If $j^{\star}\leq n/2$, then
$2\|{\sum_{i=1}^{j^{\star}}\boldsymbol{x}_{\pi_{k+1}(\pi_{k}^{-1}(i))}^{k}}\|_{2}\leq 2\max_{j\in[n]}\|{\sum_{i=1}^{j}\mathbf{1}\mathopen{}\mathclose{{}\left[{\epsilon_{i}^{k}=1}}\right]\boldsymbol{x}_{i}^{k}}\|_{2}\leq\max_{j\in[n]}\|{\sum_{i=1}^{j}\boldsymbol{x}_{i}^{k}}\|_{2}+\max_{j\in[n]}\|{\sum_{i=1}^{j}\epsilon_{i}^{k}\boldsymbol{x}_{i}^{k}}\|_{2}$ (203)
@@
$2(\|{\sum_{i=1}^{j^{\star}}\boldsymbol{x}_{\pi_{k+1}(\pi_{k}^{-1}(i))}^{k}}\|_{2}-\|{\sum_{i=1}^{n}\boldsymbol{x}_{i}^{k}}\|_{2})$ $\leq 2\|{\sum_{i>j^{\star}}\boldsymbol{x}_{\pi_{k+1}(\pi_{k}^{-1}(i))}^{k}}\|_{2}\leq 2\max_{j\in[n]}\|{\sum_{i=1}^{j}\mathbf{1}\mathopen{}\mathclose{{}\left[{\epsilon_{i}^{k}=-1}}\right]\boldsymbol{x}_{i}^{k}}\|_{2}$ (204)
$\leq\max_{j\in[n]}\|{\sum_{i=1}^{j}\boldsymbol{x}_{i}^{k}}\|_{2}+\max_{j\in[n]}\|{-\sum_{i=1}^{j}\epsilon_{i}^{k}\boldsymbol{x}_{i}^{k}}\|_{2}.$ (205)
@@
and that SGD 29 with thinned reordering (Alg. 2) satisfies the prefix discrepancy bound
$\max_{j\in[n]}\|{\sum_{i=1}^{j}\epsilon_{i}^{k}\boldsymbol{x}_{i}^{k}}\|_{2}\leq 2\tilde{A}\max_{i\in[n]}\|{\boldsymbol{x}_{i}^{k}-\bar{\boldsymbol{x}}^{k}}\|_{2}\quad\text{for}\quad\epsilon_{i}^{k}\triangleq 2(\mathbf{1}\mathopen{}\mathclose{{}\left[{\boldsymbol{x}_{i}^{k}\in\mathcal{X}_{\textup{out}}^{k}}}\right]-1),\quad\bar{\boldsymbol{x}}^{k}\triangleq\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{x}_{i}^{k},$ (207)
@@
Fix any epoch $k\in[K]$. With probability at least $1-\frac{\delta}{2}-\delta^{\prime}$, thinned reordering (Alg. 2) with LKH$(\delta)$ satisfies the prefix discrepancy bound 207 with
$\tilde{A}=\sqrt{\log(\frac{2n(\log(n/2)+1)}{\delta})\mathopen{}\mathclose{{}\left[e^{2}\,\mathrm{rank}_{\epsilon_{k}}(\mathbf{X}^{k})+e\log(\frac{n}{\delta^{\prime}})}\right]+1}$ (213)
for $\mathbf{X}^{k}\triangleq[\boldsymbol{x}_{1}^{k},\dots,\boldsymbol{x}_{n}^{k}]^{\top}$, $\epsilon_{k}\triangleq{\max_{i\in[n]}\|{\boldsymbol{x}_{i}^{k}-\bar{\boldsymbol{x}}^{k}}\|_{2}}/{\sqrt{n}},$ and $\bar{\boldsymbol{x}}^{k}\triangleq\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{x}_{i}^{k}.$
@@
$\|{\sum_{i=1}^{j}\epsilon_{i}^{k}\boldsymbol{x}_{i}^{k}}\|_{2}=\|{\sum_{i=1}^{j}\boldsymbol{x}_{i}^{k}-\sum_{i=1}^{j}\mathbf{1}\mathopen{}\mathclose{{}\left[{\boldsymbol{x}_{i}^{k}\in\mathcal{X}_{\textup{out},j}^{k}}}\right]\boldsymbol{x}_{i}^{k}}\|_{2}=2j\|{(\mathbf{X}^{k})^{\top}(\boldsymbol{p}_{\textup{in}}^{j}-\boldsymbol{p}_{\textup{out}}^{j})}\|_{2}=2j\operatorname{MMD}_{\mathbf{X}^{k}(\mathbf{X}^{k})^{\top}}(\boldsymbol{p}_{\textup{in}}^{j},\boldsymbol{p}_{\textup{out}}^{j})$ (214)
@@
Since LKH$(\delta)$ is an online algorithm that assigns signs $(\epsilon_{i}^{k},\epsilon_{i+1}^{k}=1-\epsilon_{i}^{k})$ to the points $(\boldsymbol{x}_{i}^{k},\boldsymbol{x}_{i+1}^{k})$ sequentially, we can view $\mathcal{X}_{\textup{out},j}^{k}$ as the output of LKH$(\delta)$ applied to $\mathcal{X}_{\textup{in},j}^{k}$ with $n_{\textup{out}}=\frac{j}{2}$ and the linear kernel $\mathbf{k}(\boldsymbol{x},\boldsymbol{y})=\langle{\boldsymbol{x}},{\boldsymbol{y}}\rangle$ for each $j\in[n]$. Therefore, we may invoke the established LKH$(\delta)$ sub-Gaussian constants $\nu_{j}$ of Prop. B.3, Thm. 1, the union bound, and the definition of $\epsilon$-rank (Def. 4) to deduce that
$\max_{j\in[n]}\|{\sum_{i=1}^{j}\epsilon_{i}^{k}\boldsymbol{x}_{i}^{k}}\|_{2}^{2}$ $\leq\max_{j\in[n]}4j^{2}\nu_{j}^{2}\mathopen{}\mathclose{{}\left[e^{2}r+e\log(\frac{n}{\delta^{\prime}})}\right]+\sigma_{r+1}(\mathbf{X}^{k})^{2}\frac{4j^{2}}{j}\leq c^{2}\tilde{A}^{2}$ (215)

### v7

@@ C.2 Proof of kernel max seminorm bound 13
The triangle inequality and the covering property 133 together imply that, with probability 111,
@@
for each $f\in\mathcal{F}$. Since $s\mapsto e^{ts}$ is increasing, the bound 138, the assumed sub-Gaussianity (Def. A.2), and the fact that $\mathcal{C}_{\epsilon,\mathcal{F}}$ belongs to $\mathbb{B}_{\mathcal{F}}$ imply that
@@ C.3 Proof of Lipschitz kernel max seminorm bound 15
Since $\mathbb{P}(\mathcal{E}^{c})\leq\delta/2$, it only remains to upper bound $\mathopen{}\mathclose{{\left\|{(\mathbb{P}_{\textup{in}}-\mathbb{P}_{\textup{out}})\mathbf{k}}}}\right\|_{\mathcal{Z},\mathcal{Z}}$ on $\mathcal{E}$ with probability at least $1-\delta^{\prime}/2$.
@@
Suppose $(X_{\theta})_{\theta\in\Theta}$ is a sub-Gaussian process with respect to $\rho$ on an event $\mathcal{E}$, and define the diameter $\mathop{\mathrm{diam}}(\Theta,\rho)\triangleq\sup_{\theta,\theta^{\prime}\in\Theta}\rho(\theta,\theta^{\prime})$, the covering number
@@
$\mathbb{P}_{\mathcal{E}}(\sup_{\theta,\theta^{\prime}\in\Theta}|X_{\theta}-X_{\theta^{\prime}}|\geq 8(\mathcal{J}(\Theta,\rho)+t))\leq 2\exp(-t^{2}/\mathop{\mathrm{diam}}(\Theta,\rho)^{2})\quad\text{for all}\quad t>0.$ (151)
@@
Consider the bounded-Hölder metric $\rho$ 149 for a kernel $\mathbf{k}$ on $\mathcal{X}=\mathbb{R}^{d}$ and a finite set $\mathcal{Z}\subset\mathcal{X}$. If $\mathbf{Z}$ is a matrix with one row corresponding to each element of $\mathcal{Z}$, $r=\mathrm{rank}({\mathbf{Z}})$, and $R=\max_{\boldsymbol{z}\in\mathcal{Z}}\|{\boldsymbol{z}}\|_{2}$, then, in the notation of Lem. C.5,
@@
$\mathop{\mathrm{diam}}(\mathcal{Z},\rho)$ $\leq D\triangleq\min(c,\nu\sqrt{\frac{32}{3}}\max_{\boldsymbol{z}\in\mathcal{Z}}\sqrt{\mathbf{k}(\boldsymbol{z},\boldsymbol{z})}),\quad\text{and}\quad$ (153)
@@
The diameter bound 153 follows directly from the definition of $\rho$ 149 and the fact $\max_{\boldsymbol{z},\boldsymbol{z}^{\prime}\in\mathcal{Z}}\|{\boldsymbol{z}-\boldsymbol{z}^{\prime}}\|_{2}\leq 2R$.
To establish the covering number bound 152, we let $\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$ be a compact singular value decomposition of $\mathbf{Z}$ so that
@@
so that $\mathbf{V}\mathcal{C}$ satisfies the criteria of 157. Since $|\mathbf{V}\mathcal{C}|\leq|\mathcal{C}|\leq(1+4R/\epsilon^{2})^{r}$ by Wainwright (2019, Lem. 5.2), we must also have $|\mathcal{C}_{\textup{ext}}|\leq(1+4R/\epsilon^{2})^{r}$.
Now, since $\mathcal{C}_{\textup{ext}}$ has minimum cardinality amongst sets satisfying 157, for each $\boldsymbol{z}^{\prime}\in\mathcal{C}_{\textup{ext}}$, there is some $\boldsymbol{z}\in\mathcal{Z}$ satisfying $\|{\boldsymbol{z}^{\prime}-\boldsymbol{z}}\|_{2}\leq\epsilon^{2}/2$ (or else $\boldsymbol{z}^{\prime}$ would be superfluous). Hence, there exists a set $\mathcal{C}_{\textup{int}}\subseteq\mathcal{Z}$ satisfying
@@
Hence, for $u=\frac{c\epsilon}{2\sqrt{R}}$, $\mathcal{N}(u;\mathcal{Z},\rho)\leq|\mathcal{C}_{\textup{int}}|\leq(1+c^{2}/u^{2})^{r}$. Since $\epsilon>0$ was arbitrary, we have established 152.
@@
on $\mathcal{E}$ with probability at least $1-\delta^{\prime}/2$. Combining this bound with the inequality 147 yields the result.
@@
Instantiate the notation of Algs. 1 and 172 and define the query set
@@
[…]
@@
$\leq\|{(\frac{1}{n}\mathbf{\widehat{D}})^{-1}}\|_{\max}\frac{1}{n}\|{\mathbf{A}\mathbf{1}_{n}-\widehat{\mathbf{A}}\mathbf{1}_{n}}\|_{\infty}\|{\mathbf{V}}\|_{\max}.$ (187)
@@
$\max_{j\in[n]}\|{\sum_{i=1}^{j}\boldsymbol{x}_{\pi_{k+1}(\pi_{k}^{-1}(i))}^{k}}\|_{2}\leq\frac{1}{2}\max_{j\in[n]}\|{\sum_{i=1}^{j}\boldsymbol{x}_{i}^{k}}\|_{2}+\frac{1}{2}\max_{j\in[n]}\|{\sum_{i=1}^{j}\eta_{i}^{k}\boldsymbol{x}_{i}^{k}}\|_{2}+\|{\sum_{i=1}^{n}\boldsymbol{x}_{i}^{k}}\|_{2}$ (199)
where $\pi_{k}^{-1}$ is the inverse permutation of $\pi_{k}$ and $\eta_{i}^{k}\triangleq 2(\mathbf{1}\mathopen{}\mathclose{{\left[{\boldsymbol{x}_{i}^{k}\in\mathcal{X}_{\textup{out}}^{k}}}}\right]-1)$.
@@
Fix any $j^{\star}\in\mathop{\mathrm{argmax}}_{j\in[n]}\|{\sum_{i=1}^{j}\boldsymbol{x}_{\pi_{k+1}(\pi_{k}^{-1}(i))}^{k}}\|_{2}$. If $j^{\star}\leq n/2$, then
$2\|{\sum_{i=1}^{j^{\star}}\boldsymbol{x}_{\pi_{k+1}(\pi_{k}^{-1}(i))}^{k}}\|_{2}\leq 2\max_{j\in[n]}\|{\sum_{i=1}^{j}\mathbf{1}\mathopen{}\mathclose{{\left[{\eta_{i}^{k}=1}}}\right]\boldsymbol{x}_{i}^{k}}\|_{2}\leq\max_{j\in[n]}\|{\sum_{i=1}^{j}\boldsymbol{x}_{i}^{k}}\|_{2}+\max_{j\in[n]}\|{\sum_{i=1}^{j}\eta_{i}^{k}\boldsymbol{x}_{i}^{k}}\|_{2}$ (200)
@@
$2(\|{\sum_{i=1}^{j^{\star}}\boldsymbol{x}_{\pi_{k+1}(\pi_{k}^{-1}(i))}^{k}}\|_{2}-\|{\sum_{i=1}^{n}\boldsymbol{x}_{i}^{k}}\|_{2})$ $\leq 2\|{\sum_{i>j^{\star}}\boldsymbol{x}_{\pi_{k+1}(\pi_{k}^{-1}(i))}^{k}}\|_{2}\leq 2\max_{j\in[n]}\|{\sum_{i=1}^{j}\mathbf{1}\mathopen{}\mathclose{{\left[{\eta_{i}^{k}=-1}}}\right]\boldsymbol{x}_{i}^{k}}\|_{2}$ (201)
$\leq\max_{j\in[n]}\|{\sum_{i=1}^{j}\boldsymbol{x}_{i}^{k}}\|_{2}+\max_{j\in[n]}\|{-\sum_{i=1}^{j}\eta_{i}^{k}\boldsymbol{x}_{i}^{k}}\|_{2}.$ (202)
@@
and that SGD 26 with thinned reordering (Alg. 2) satisfies the prefix discrepancy bound
$\max_{j\in[n]}\|{\sum_{i=1}^{j}\eta_{i}^{k}\boldsymbol{x}_{i}^{k}}\|_{2}\leq 2\tilde{A}\max_{i\in[n]}\|{\boldsymbol{x}_{i}^{k}-\bar{\boldsymbol{x}}^{k}}\|_{2}\quad\text{for}\quad\eta_{i}^{k}\triangleq 2(\mathbf{1}\mathopen{}\mathclose{{\left[{\boldsymbol{x}_{i}^{k}\in\mathcal{X}_{\textup{out}}^{k}}}}\right]-1),\quad\bar{\boldsymbol{x}}^{k}\triangleq\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{x}_{i}^{k},$ (204)
@@
Fix any epoch $k\in[K]$. With probability at least $1-\frac{\delta}{2}-\delta^{\prime}$, thinned reordering (Alg. 2) with LKH$(\delta)$ satisfies the prefix discrepancy bound 204 with
$\tilde{A}=\sqrt{\log(\frac{2n(\log(n/2)+1)}{\delta})\mathopen{}\mathclose{{\left[e^{2}\,\mathrm{rank}_{\epsilon_{k}}(\mathbf{X}^{k})+10e\log(\frac{n}{\delta^{\prime}})}}\right]}$ (210)
for $\mathbf{X}^{k}\triangleq[\boldsymbol{x}_{1}^{k},\dots,\boldsymbol{x}_{n}^{k}]^{\top}$, $\epsilon_{k}\triangleq\max_{i\in[n]}\ \textstyle\frac{\sqrt{9e\log(2n\log(en/2)/\delta)\log(n/\delta^{\prime})}\|{\boldsymbol{x}_{i}^{k}-\bar{\boldsymbol{x}}^{k}}\|_{2}}{\sqrt{n}},$ and $\bar{\boldsymbol{x}}^{k}\triangleq\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{x}_{i}^{k}.$
@@
$\|{\sum_{i=1}^{j}\eta_{i}^{k}\boldsymbol{x}_{i}^{k}}\|_{2}=\|{\sum_{i=1}^{j}\boldsymbol{x}_{i}^{k}-\sum_{i=1}^{j}\mathbf{1}\mathopen{}\mathclose{{\left[{\boldsymbol{x}_{i}^{k}\in\mathcal{X}_{\textup{out},j}^{k}}}}\right]\boldsymbol{x}_{i}^{k}}\|_{2}=2j\|{(\mathbf{X}^{k})^{\top}(\boldsymbol{p}_{\textup{in}}^{j}-\boldsymbol{p}_{\textup{out}}^{j})}\|_{2}=2j\operatorname{MMD}_{\mathbf{X}^{k}(\mathbf{X}^{k})^{\top}}(\boldsymbol{p}_{\textup{in}}^{j},\boldsymbol{p}_{\textup{out}}^{j})$ (211)
@@
Since LKH$(\delta)$ is an online algorithm that assigns signs $(\eta_{i}^{k},\eta_{i+1}^{k}=1-\eta_{i}^{k})$ to the points $(\boldsymbol{x}_{i}^{k},\boldsymbol{x}_{i+1}^{k})$ sequentially, we can view $\mathcal{X}_{\textup{out},j}^{k}$ as the output of LKH$(\delta)$ applied to $\mathcal{X}_{\textup{in},j}^{k}$ with $n_{\textup{out}}=\frac{j}{2}$ and the linear kernel $\mathbf{k}(\boldsymbol{x},\boldsymbol{y})=\langle{\boldsymbol{x}},{\boldsymbol{y}}\rangle$ for each $j\in[n]$. Therefore, we may invoke the established LKH$(\delta)$ sub-Gaussian constants $\nu_{j}$ of Prop. B.3, Thm. 1, the union bound, and the definition of $\epsilon$-rank (Def. 4) to deduce that
$\max_{j\in[n]}\|{\sum_{i=1}^{j}\eta_{i}^{k}\boldsymbol{x}_{i}^{k}}\|_{2}^{2}$ $\leq\max_{j\in[n]}4j^{2}\nu_{j}^{2}\mathopen{}\mathclose{{\left[e^{2}r+e\log(\frac{n}{\delta^{\prime}})}}\right]+\sigma_{r+1}(\mathbf{X}^{k})^{2}\frac{4j^{2}}{j}\leq c^{2}\tilde{A}^{2}$ (212)